A given mass of a gas occupies 150cm cube at 20 degrees Celsius and pressure at 1.013*(10*5) Newton per meter. Calculate the temperature at which the volume is doubled at the same pressure?

1
Expert's answer
2022-02-16T05:06:01-0500

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


"P_1=P_2"


"\frac{150}{273.15+20}=\frac{2(150)}{T}"


"T=586.3K"